# Strong anharmonicity dictates ultralow thermal conductivities of type-I clathrates

### v1

@@ Abstract
Type-I clathrate solids have attracted significant interest due to their ultralow thermal conductivities and subsequent promise for thermoelectric applications, yet the mechanisms underlying these properties are not well understood. Here, we extend the framework of vibrational dynamical mean-field theory (VDMFT) to calculate temperature-dependent thermal transport properties of $X_{8}$a16Ge30, where $X=$ Ba, Sr, using a many-body Green’s function approach. We find that nonresonant scattering between cage acoustic modes and rattling modes leads to a reduction of acoustic phonon lifetimes and thus thermal conductivities. Moreover, we find that the moderate temperature dependence of conductivities above 300 K, which is consistent with experimental measurements, cannot be reproduced by standard perturbation theory calculations, which predict a $T^{-1}$ dependence. Therefore, we conclude that nonperturbative anharmonic effects, including four- and higher-phonon scattering processes, are responsible for the ultralow thermal conductivities of type-I clathrates.
@@ I Introduction
Good thermoelectric materials have high electronic conductivities but low thermal conductivities, which presents a challenging problem in materials design [1]. One promising class of materials are type-I clathrate solids, which contain cage-like chemical structures of covalently bonded atoms that host loosely bound “guest” atoms (Fig. 1a), although the mechanism responsible for their low thermal conductivity has been debated [2, 3, 4, 5, 6, 7]. The rattling displacements of the guest atoms generate dispersionless, low-frequency optical phonon branches that intersect the acoustic phonon branches of the cage lattice [3]. The hybridization of these branches suppresses the group velocity of the acoustic phonons, which also inherit the reduced lifetimes of the strongly anharmonic optical rattling phonons [4, 8]. More exotic behaviors have also been proposed, including a phonon Kondo effect [6].
A standard approach to calculate the thermal conductivity tensor $\kappa^{ij}$ is based on the Boltzmann transport equation in the single-mode approximation (BTE-SMA) [9],
@@
where $V$ is the cell volume, $\lambda$ is an index over phonon modes in the system, $C_{\lambda}(\bm{k})$ is the heat capacity of mode $\lambda$, and $v^{i}_{\lambda}(\bm{k})$ is the phonon group velocity of mode $\lambda$ in Cartesian direction $i$. The mode lifetimes $\tau_{\lambda}(\bm{k})$ are commonly obtained by treating three-phonon processes to second order in perturbation theory (PT). Given the strong anharmonicity required to suppress the lifetimes of acoustic phonons in materials with low thermal conductivities, it is natural to question the quantitative accuracy of this approach. In particular, at temperatures above the Debye temperature, the heat capacity is constant, and PT predicts that the lifetimes decrease as $\tau\propto T^{-1}$, such that $\kappa\propto T^{-1}$. Experimentally observed deviations from this temperature dependence are indicative of stronger anharmonicity [10, 11], and their theoretical explanation requires more advanced methods. Approaches relying on equilibrium and nonequilibrium molecular dynamics (MD) simulations [2, 12, 13, 14] include nonperturbative effects of strong anharmonic interactions, but the computational expense associated with such direct simulation, especially when considering finite-size effects, makes them intractable for systems with complex unit cells. Furthermore, obtaining systematic or mechanistic insight from such large-scale atomistic simulations can be challenging relative to approaches, such as the BTE-SMA, which calculate phonon-mode resolved thermal conductivities.
Recently, we introduced vibrational dynamical mean-field theory (VDMFT) as an affordable but nonperturbative computational method for simulating anharmonic lattice dynamics [15], and we used it to calculate the phonon spectral function of the type-I clathrates Ba8Ga16Ge30 (BaGG) and Sr8Ga16Ge30 (SrGG) [8]. Here, we extend the VDMFT methodology to enable the calculation of conductivities (neglecting vertex corrections) [16], which we then use to calculate the temperature-dependent conductivities of the same clathrate systems studied in our previous work. We also briefly evaluate the evidence for a phonon Kondo effect [6], which our method is uniquely suited to study, given the success of DMFT and its associated impurity problem in explaining Kondo physics in electronic systems.
@@
Through our VDMFT calculations, we confirm that anharmonic interactions between acoustic modes of the cage lattice and flat optical modes corresponding to rattling motions of guest atoms inside 24-atom cages are responsible for the ultralow thermal conductivities of these materials. Furthermore, through a nonperturbative description of this anharmonicity, we find that the thermal conductivity has a relatively moderate dependence on the temperature above 300 K, in agreement with experimental observations. These results cannot be captured, even qualitatively, using conventional lowest-order PT, suggesting that nonperturbative effects of anharmonicity, including four-phonon and higher-order scattering processes, are key to understanding the reduced thermal transport of type-I clathrates.
@@ II.1 Clathrate model
To study the vibrational structure and thermal transport properties of type-I clathrates, we use a coarse-grained model of the material [8], illustrated in Fig. 1b. We model dodecahedral and tetrakaidecahedral cages with single “hollow” atoms, which are arranged on an FCC lattice and interact through a Lennard-Jones (LJ) potential. Guest atoms are described by smaller atoms at each FCC site, which interact with the cage atoms at those sites through anharmonic, quartic potentials. The Hamiltonian for our clathrate model is given by
@@
where the primed summation indicates that only cages and guests on the same lattice site interact. Here, $\bm{m},\bm{n}$ are lattice translation vectors, $\alpha,\beta$ are indices over atoms in the unit cell, and $i$ is an index over the Cartesian directions. The position, momentum, and mass of atom $\alpha$ in cell $\bm{m}$ are given by $\bm{r}_{\bm{m}\alpha}$, $\bm{p}_{\bm{m}\alpha}$, and $m_{\alpha}$, respectively. Further details about our model, including parameters for the LJ and quartic potentials, are given in our previous work [8]. Importantly, the symmetry of our model implies that cubic force constants involving guest atoms are identically zero, i.e., $\partial^{3}\mathcal{V}/\partial u_{\bm{m}\alpha i}\partial u_{\bm{n}\beta j}\partial u_{\bm{l}\gamma k}=0$, where $\alpha,\beta$ and/or $\gamma$ are guest atoms, although this is not true of the fully atomic material. This difference has important implications for the performance of perturbation theory and comparison to previous works, as discussed below.
@@ II.2 Anharmonic lattice dynamics with vibrational dynamical mean-field theory
expressed in the basis of of translational symmetry adapted, mass-weighted atomic displacements of atom $\alpha$ in Cartesian direction $i$,
@@
In VDMFT, the dynamics of the periodic lattice are mapped onto an “impurity problem” that consists of a single unit cell coupled to a fictitious bath of harmonic oscillators through a tailored spectral density [24, 25, 26, 27, 15, 8]. The local self-energy in a single unit cell, $\bm{\pi}(\omega)$, is calculated through the exact solution of the impurity problem and is used to approximate the self-energy of the lattice GF, $\pi_{\alpha i,\beta j}(\bm{k},\omega)\approx\pi_{\alpha i,\beta j}(\omega)$, thus providing a nonperturbative treatment of anharmonicity. The lattice GF and impurity problem are iteratively updated until self-consistency is achieved. We perform SCP calculations using the ALAMODE package [28] (details given in Supplemental Material, SM) and use that quasiparticle basis in combination with VDMFT, following the approach described in detail in our previous work [15, 8]. As in our previous work, the impurity GF is calculated by neglecting nuclear quantum effects and treating the nuclei classically.
@@
We assess the accuracy of our SCP+VDMFT approach by comparing the spectral function of SrGG, the most anharmonic system studied here, with the exact spectral function computed using molecular dynamics (MD) simulations of a large supercell with periodic boundary conditions. The spectral functions at the $X$ point of the BZ are illustrated in Fig. 1d along with the frequency shifts and linewidths calculated on a 4$\times$4$\times$4 grid of the BZ, demonstrating excellent agreement between SCP+VDMFT and MD. We also compare linewidths to those calculated with lowest-order perturbation theory (PT) of three-phonon scattering processes using the SCP quasiparticle basis, which is the conventional approach for computing phonon linewidths [5]. Details regarding the MD and SCP+PT calculations are given in the SM. Unlike MD and SCP+VDMFT, Fig. 1d shows that SCP+PT inaccurately predicts negligible broadening for all modes in SrGG. This failure of SCP+PT has significant implications for calculations of the thermal conductivity, which rely on the accurate description of phonon linewidths, as described below.
@@ II.3.1 Green-Kubo formalism
While the response function, $\chi_{ij}(t)$, depends on products of four bosonic operators, or a two-particle GF, it can be approximated as a product of two single-particle GFs by neglecting vertex corrections [31]. This approximation, which is exact in infinite dimensions [24], is expected to be valid for materials with low thermal conductivity, in which resistive, Umklapp scattering processes are dominant over momentum-conserving, normal scattering processes that are characteristic of the hydrodynamic regime of thermal transport [32]. This approximation yields the thermal conductivity [16],
@@
where $n_{T}(\omega)=\big{(}e^{\hbar\omega/k_{B}T}-1\big{)}^{-1}$ is the Bose-Einstein distribution. Equation (14) contains both intraband ($\lambda=\lambda^{\prime}$) and interband ($\lambda\neq\lambda^{\prime}$) contributions [33, 34, 16]. The interband contributions become important in materials for which phonon broadening is on the same order as energy spacing between phonon bands, as seen in the clathrates studied here, especially SrGG (Fig. 1c).
@@ II.3.2 Boltzmann transport equation within the single mode approximation
The full expression for the thermal conductivity given by Eq. (14) can be simplified under the assumption that the phonon quasiparticle picture is valid such that interband transitions are negligible and that the phonon spectral densities have Lorenztian lineshapes centered at the harmonic phonon frequencies $\omega_{\lambda}(\bm{k})$ with widths $\Gamma_{\lambda}(\bm{k})$. These approximations yield an expression for the thermal conductivity that is identical to the BTE-SMA expression in Eq. (1), with $v_{\lambda}^{i}(\bm{k})=v_{\lambda,\lambda}^{i}(\bm{k})$ and $\tau_{\lambda}(\bm{k})=1/2\Gamma_{\lambda}(\bm{k})$.
@@ III.1 Intraband transport
The SM shows the calculated mean-free-paths of acoustic modes of GG, BaGG, and SrGG at various temperatures. At all temperatures, the filled clathrates have mean-free-paths that are an order of magnitude or more smaller than those of the empty clathrate. This reduction is due to a sharp decrease in phonon lifetimes for the filled clathrate acoustic modes in a wide frequency range, confirming the role of nonresonant scattering in these materials [4]. In contrast, the acoustic phonon mode velocities are unchanged near the BZ center, and they only decrease as they approach the zone edge and gain significant character of the X(2) guest atoms.
@@ III.2 Interband transport
Due to the strong anharmonicity and significant mode mixing of the filled clathrates, especially near the avoided crossing of the cage-acoustic and guest-optical modes, we investigate the role of interband transport in the overall thermal conductivity [33, 35]. Figure 3a shows the percent of the total thermal conductivity that is due to interband transport. As expected, GG has negligible interband transport at all temperatures because of its relatively harmonic lattice dynamics, for which the particle-like picture of thermal transport is dominant. The filled clathrates, however, have significant contributions from interband transport that increases with increasing temperature. At 600 K, 8% and 11% of the total thermal conductivity for BaGG and SrGG, respectively, are due to interband transport.
@@ III.3 The role of nonperturbative effects on thermal conductivity
Figure 4: (a) Thermal conductivities of GG (blue), BaGG (green), and SrGG (red) calculated using the Green-Kubo formalism [Eq. (14)] with spectral functions calculated using SCP+VDMFT (triangles) and using the BTE formalism [Eq. (1)] with lifetimes calculated using SCP+PT (circles). All calculations are sampled on a $12\times 12\times 12$ Gamma-centered grid of the BZ. (b) Thermal conductivities at higher temperatures, where black points are experimental measurements from Ref. [11]. Grey lines indicating different power laws are a guide to the eye.
@@
To better understand the role of nonperturbative effects on the thermal conductivity, we compare the thermal conductivity across materials and temperatures computed using two different methods: the Green-Kubo formalism [Eq. (14)] with spectral functions calculated using SCP+VDMFT (GK-VDMFT) and the BTE-SMA approach [Eq. (1)] with SCP phonons and lifetimes calculated using conventional lowest-order PT of the three-phonon scattering processes (BTE-PT). Details regarding the BTE-PT calculations are provided in the SM.
@@
When calculated using GK-VDMFT, the thermal conductivity for all materials peaks around 25-50 K and then decreases with increasing temperature. As discussed above, the introduction of guest atoms decreases the thermal conductivity by more than one order of magnitude, and BaGG, which features weaker cage-guest anharmonicity, has a larger thermal conductivity than SrGG at all temperatures. The BTE-PT approach results in thermal conductivity values that are consistent with GK-VDMFT for the quasi-harmonic GG, but they disagree significantly for the filled clathrates. BTE-PT predicts only a factor of two reduction between GG and BaGG, and it incorrectly predicts that the more anharmonic SrGG has a larger thermal conductivity than BaGG. These discrepancies are a result of the failure of SCP+PT to capture the short lifetimes of the filled clathrates (Fig. 1d), causing a significant overestimation of their thermal conductivities. As shown in the SM, using the BTE-SMA formalism with lifetimes determined from SCP+VDMFT calculations yields values that closely align with those calculated by the GK-VDMFT approach, suggesting that the success of the BTE-SMA depends strongly on the quality of the phonon lifetimes used and, in this case, is not due to the exclusion of interband transport. The SM also illustrates comparisons to thermal conductivities calculated according to the classical analog of Eq. (10) using molecular dynamics simulations, showing consistency with the GK-VDMFT results. These comparisons suggest that the ultralow thermal conductivities of the filled clathrates result from four-phonon and higher-order scattering processes that are not described by SCP+PT but that are inherently described by our SCP+VDMFT approach, highlighting the importance of nonperturbative anharmonic effects on observables like the thermal conductivity. In Fig. 4b, we compare our calculated values of the thermal conductivity for BaGG to experimental data, focusing on high-temperature data, where intrinsic effects of anharmonicity dominate over effects due to grain boundary scattering, dynamical disorder, and nuclear quantum effects. We see excellent agreement, validating our approach. At lower temperatures, phonon-grain boundary scattering can become important [4], as discussed in the SM.
@@
Finally, we consider the temperature dependence of the total thermal conductivity (Fig. 4b), as strong interactions can produce anomalous behavior, as in electronic transport properties [36]. Here, we focus on the high-temperature dependence, as our calculations neglect nuclear quantum effects and grain boundary scattering, which may affect the behavior of thermal conductivities at low temperatures. For all materials, BTE-PT predicts a $T^{-1}$ temperature dependence of the thermal conductivity. This temperature dependence is dictated by that of the phonon lifetimes, for which a $T^{-1}$ scaling is inherent to the PT formalism at temperatures greater than the phonon frequency (see SM). In contrast, the GK-VDMFT approach predicts different high-temperature behavior for the different clathrates. BaGG has an intermediate scaling of $\sim T^{-0.9}$, and SrGG, the most anharmonic system, has the most moderate scaling, $\sim T^{-0.6}$.
@@
The high-temperature dependence of the thermal conductivity primarily arises from that of the acoustic phonon lifetimes. Thus, we analyze the temperature dependence of longitudinal and transverse acoustic phonon modes near the BZ center of BaGG and SrGG, as illustrated in Fig. 5. While the mode lifetimes of BaGG largely follow a $T^{-1}$ behavior, the mode lifetimes of SrGG deviate significantly from the $T^{-1}$ behavior that is predicted by perturbation theory and begin to saturate to a milder scaling at 50 K. This high-temperature saturation of the phonon lifetimes, and thus thermal conductivities, is reminiscent of “resistivity saturation” in metals [36, 37, 38, 39] and is a hallmark of nonperturbative interactions. This result indicates that the acoustic modes of the SrGG cage framework are subject to multiphonon scattering processes that are not captured by conventional perturbation theory. Despite their relatively narrow linewidths and long lifetimes (Fig. 1), these acoustic phonons’ low energies and strong coupling to Sr(2) rattling modes makes them more likely to participate in multiphonon scattering processes. Thus, we find that including nonperturbative descriptions of anharmonicity is essential in accurately capturing thermal transport properties, especially at higher temperatures where multiphonon processes are activated.

### v2

@@ Abstract
Type-I clathrate solids have attracted significant interest due to their ultralow thermal conductivities and subsequent promise for thermoelectric applications, yet the mechanisms underlying these properties are not well understood. Here, we extend the framework of vibrational dynamical mean-field theory (VDMFT) to calculate temperature-dependent thermal transport properties of solids using a many-body Green’s function approach. When applied to a coarse-grained description of $X_{8}$a16Ge30, where $X=$ Ba, Sr, we find that nonresonant scattering between cage acoustic modes and rattling modes leads to a reduction of acoustic phonon lifetimes and thus thermal conductivities. Moreover, we find that the moderate temperature dependence of conductivities above 300 K, which is consistent with experimental measurements, cannot be reproduced by textbook perturbation theory calculations, which predict a $T^{-1}$ dependence. Therefore, we suggest that nonperturbative anharmonic effects, including four- and higher-phonon scattering processes, are responsible for the ultralow thermal conductivities of type-I clathrates.
@@ I Introduction
Good thermoelectric materials, which have the potential to convert waste heat to useful electricity [1, 2, 3], have high electronic conductivities but low thermal conductivities, which presents a challenging problem in materials design [4]. One promising class of materials are type-I clathrate solids, which contain cage-like chemical structures of covalently bonded atoms that host loosely bound “guest” atoms (Fig. 1a), although the mechanism responsible for their low thermal conductivity has been debated [5, 6, 7, 8, 9, 10, 11]. The rattling displacements of the guest atoms generate dispersionless, low-frequency optical phonon branches that intersect the acoustic phonon branches of the cage lattice [6]. The hybridization of these branches suppresses the group velocity of the acoustic phonons, which also inherit the reduced lifetimes of the strongly anharmonic optical rattling phonons [7, 12]. More exotic behaviors have also been proposed, including a phonon Kondo effect [9].
One of the simplest microscopic approaches to calculate the thermal conductivity tensor, $\kappa^{ij}$, is based on the Boltzmann transport equation in the single-mode approximation (BTE-SMA) [13],
@@
where $V$ is the cell volume, $\lambda$ is an index over phonon modes in the system, $C_{\lambda}(\bm{k})$ is the heat capacity of mode $\lambda$, and $v^{i}_{\lambda}(\bm{k})$ is the phonon group velocity of mode $\lambda$ in Cartesian direction $i$. The mode lifetimes, $\tau_{\lambda}(\bm{k})$, are commonly obtained by treating three-phonon processes to second order in perturbation theory (PT). Given the strong anharmonicity required to suppress the lifetimes of acoustic phonons in materials with low thermal conductivities, it is natural to question the quantitative accuracy of this approach. In particular, at temperatures above the Debye temperature, the heat capacity is constant, and PT predicts that the lifetimes decrease as $\tau\propto T^{-1}$, such that $\kappa\propto T^{-1}$. Experimentally observed deviations from this temperature dependence are indicative of stronger anharmonicity [14, 15], and their theoretical explanation requires more advanced methods. For example, four-phonon scattering and temperature-dependent phonon renormalization can be accounted for using higher-order PT [16, 17, 18] or self-consistent phonon theory [19, 8, 18]. Additionally, recent approaches based on Wigner transport and Green-Kubo linear response theory [20, 21, 22] are able to describe effects, such as interband transport (further discussed in Sec. II.C), that are neglected in the BTE-SMA. These methods have been shown to successfully capture deviations from the $T^{-1}$ behavior in various materials with low thermal conductivities.
Nonperturbative methods are not reliant on finite-order PT and therefore inherently include all orders of phonon scattering processes. Approaches based on equilibrium and nonequilibrium molecular dynamics (MD) simulations [5, 23, 24, 25, 26] include nonperturbative effects of strong anharmonic interactions, but the computational expense associated with such direct simulation, especially when considering finite-size effects, makes them intractable for systems with complex unit cells. Furthermore, obtaining systematic or mechanistic insight from such large-scale atomistic simulations can be challenging relative to approaches, such as the BTE-SMA, which calculate phonon-mode resolved thermal conductivities. Recently, we introduced vibrational dynamical mean-field theory (VDMFT) as an affordable but nonperturbative computational method for simulating anharmonic lattice dynamics [27], and we used it to calculate the phonon spectral function of the type-I clathrates Ba8Ga16Ge30 (BaGG) and Sr8Ga16Ge30 (SrGG) [12].
The present work has two goals. The first goal, which is methodological, is to extend VDMFT to enable the calculation of thermal conductivities. Because DMFT is a theory of the one-body Green’s function, it does not formally predict two-body Green’s functions needed to calculate transport coefficients. In electronic structure theory, where DMFT is a mature method, this limitation is commonly addressed by approximating the two-body Green’s function in terms of the one-body Green’s function (i.e., neglecting vertex corrections [28, 29]), yielding electronic resistivities that exhibit signatures of strong electron correlations and good agreement with experiments [30, 31, 32]. Here, we provide the first implementation and test of such a framework for lattice thermal conductivities. We choose to test these methodological developments on model systems that are simple enough to be studied in detail with the present theory but complex enough to be representative of real materials of interest. This defines our second goal, which is to provide insight into the possible roles of strong anharmonicity on the temperature-dependent transport properties of type-I clathrate materials.
@@
Through our VDMFT calculations on model clathrate systems, we confirm that anharmonic interactions between acoustic modes of the cage lattice and flat optical modes corresponding to rattling motions of guest atoms inside 24-atom cages are responsible for the ultralow thermal conductivities of these materials. Furthermore, through a nonperturbative description of this anharmonicity, we find that the thermal conductivity has a relatively moderate dependence on the temperature above 300 K, consistent with experimental observations. Within our model, these results cannot be captured, even qualitatively, using conventional lowest-order PT, suggesting that nonperturbative effects of anharmonicity, including four-phonon and higher-order scattering processes, are important for understanding the reduced thermal transport of type-I clathrates. We also briefly evaluate the evidence for a phonon Kondo effect [9], which our method is uniquely suited to study, given the success of DMFT and its associated impurity problem in explaining Kondo physics in electronic systems.
@@ II.1 Clathrate model
To study the vibrational structure and thermal transport properties of type-I clathrates, we use a coarse-grained model of the material [12], illustrated in Fig. 1b. We model dodecahedral and tetrakaidecahedral cages with single “hollow” atoms, which are arranged on an FCC lattice and interact through a Lennard-Jones (LJ) potential. While the model neglects high-frequency intracage dynamics, the resulting modes have small group velocities and high frequencies and will therefore have minor contributions to the overall thermal conductivity [7]. Guest atoms are described by smaller atoms at each FCC site, which interact with the cage atoms at those sites through anharmonic, quartic potentials. The Hamiltonian for our clathrate model is given by
@@
where the primed summation indicates that only cages and guests on the same lattice site interact. Here, $\bm{m},\bm{n}$ are lattice translation vectors, $\alpha,\beta$ are indices over atoms in the unit cell, and $i$ is an index over the Cartesian directions. The position, momentum, and mass of atom $\alpha$ in cell $\bm{m}$ are given by $\bm{r}_{\bm{m}\alpha}$, $\bm{p}_{\bm{m}\alpha}$, and $m_{\alpha}$, respectively. Further details about our model, including parameters for the LJ and quartic potentials, are given in the Supplemental Material (SM)  [^1] See Supplemental Material at [XXX] for details about clathrate model and parameters; details of molecular dynamics, self-consistent phonon, and perturbation theory calculations; figures regarding phonon mean-free-paths and effects of grain boundary scattering; figures comparing thermal conductivities calculated using different methods; and analysis of the temperature dependence of thermal conductivities. The Supplemental Material also includes Refs. [34, 35, 36]. Although our model is parameterized against first-principles calculations [12], we note that it has vanishing cubic force constants involving guest atoms, i.e., $\partial^{3}\mathcal{V}/\partial u_{\bm{m}\alpha i}\partial u_{\bm{n}\beta j}\partial u_{\bm{l}\gamma k}=0$, where $\alpha,\beta$ and/or $\gamma$ are guest atoms, although this is not true of the fully atomistic material. This difference has important implications for the performance of PT and comparison to previous works, as discussed below.
@@ II.2 Anharmonic lattice dynamics with vibrational dynamical mean-field theory
expressed in the basis of translational symmetry adapted, mass-weighted atomic displacements of atom $\alpha$ in Cartesian direction $i$,
@@
In VDMFT, the dynamics of the periodic lattice are mapped onto an “impurity problem” that consists of a single unit cell coupled to a fictitious bath of harmonic oscillators through a tailored spectral density [44, 28, 45, 46, 27, 12]. The local self-energy in a single unit cell, $\bm{\pi}(\omega)$, is calculated through the exact solution of the impurity problem and is used to approximate the self-energy of the lattice GF, $\pi_{\alpha i,\beta j}(\bm{k},\omega)\approx\pi_{\alpha i,\beta j}(\omega)$, thus providing a nonperturbative treatment of anharmonicity. The lattice GF and impurity problem are iteratively updated until self-consistency is achieved. We perform SCP calculations using the ALAMODE package [47] (details given in the SM [33]) and use that quasiparticle basis in combination with VDMFT, following the approach described in detail in our previous work [27, 12]. As in our previous work, the impurity GF is calculated by neglecting nuclear quantum effects and treating the nuclei classically.
@@
We assess the accuracy of our SCP+VDMFT approach by comparing the spectral function of SrGG, the most anharmonic system studied here, with the exact spectral function computed using molecular dynamics (MD) simulations of a large supercell with periodic boundary conditions. The spectral functions at the $X$ point of the BZ are illustrated in Fig. 1d along with the frequency shifts and linewidths calculated on a 4$\times$4$\times$4 grid of the BZ, demonstrating excellent agreement between SCP+VDMFT and MD. We also compare linewidths to those calculated with lowest-order perturbation theory (PT) of three-phonon scattering processes using the SCP quasiparticle basis, which is the conventional approach for computing phonon linewidths [8]. Details regarding the MD and SCP+PT calculations are given in the SM [33]. Unlike MD and SCP+VDMFT, Fig. 1d shows that SCP+PT inaccurately predicts negligible broadening for all modes in SrGG. This failure of SCP+PT has significant implications for calculations of the thermal conductivity, which rely on the accurate description of phonon linewidths, as described below.
@@ II.3.1 Green-Kubo formalism
While the response function, $\chi_{ij}(t)$, depends on products of four bosonic operators, or a two-particle GF, it can be approximated as a product of two single-particle GFs by neglecting vertex corrections [51]. This approximation, which is exact in infinite dimensions [44], is expected to be valid for materials with low thermal conductivity, in which resistive, Umklapp scattering processes are dominant over momentum-conserving, normal scattering processes that are characteristic of the hydrodynamic regime of thermal transport [52]. This approximation yields the thermal conductivity [53, 54, 22, 50],
@@
where $n_{T}(\omega)=\big{(}e^{\hbar\omega/k_{B}T}-1\big{)}^{-1}$ is the Bose-Einstein distribution. Equation (14) contains both intraband ($\lambda=\lambda^{\prime}$) and interband ($\lambda\neq\lambda^{\prime}$) contributions [20, 22, 21, 50]. The interband contributions become important in materials for which phonon broadening is on the same order as energy spacing between phonon bands, as seen in the clathrates studied here, especially SrGG (Fig. 1c).
@@ II.3.2 Boltzmann transport equation within the single mode approximation
The full expression for the thermal conductivity given by Eq. (14) can be simplified under the assumption that the phonon quasiparticle picture is valid such that interband transitions are negligible and that the phonon spectral densities have Lorenztian lineshapes centered at the harmonic phonon frequencies, $\omega_{\lambda}(\bm{k})$, with widths $\Gamma_{\lambda}(\bm{k})$. These approximations yield an expression for the thermal conductivity that is identical to the BTE-SMA expression in Eq. (1), with $v_{\lambda}^{i}(\bm{k})=v_{\lambda,\lambda}^{i}(\bm{k})$ and $\tau_{\lambda}(\bm{k})=1/2\Gamma_{\lambda}(\bm{k})$.
@@ III.1 Intraband transport
The SM [33] shows the calculated mean-free-paths of acoustic modes of GG, BaGG, and SrGG at various temperatures. At all temperatures, the filled clathrates have mean-free-paths that are an order of magnitude or more smaller than those of the empty clathrate. This reduction is due to a sharp decrease in phonon lifetimes for the filled clathrate acoustic modes in a wide frequency range, confirming the role of nonresonant scattering in these materials [7]. In contrast, the acoustic phonon mode velocities are unchanged near the BZ center, and they only decrease as they approach the zone edge and gain significant character of the X(2) guest atoms.
@@ III.2 Interband transport
Due to the strong anharmonicity and significant mode mixing of the filled clathrates, especially near the avoided crossing of the cage-acoustic and guest-optical modes, we investigate the role of interband transport in the overall thermal conductivity [20, 55]. Figure 3a shows the percent of the total thermal conductivity that is due to interband transport. As expected, GG has negligible interband transport at all temperatures because of its relatively harmonic lattice dynamics, for which the particle-like picture of thermal transport is dominant. The filled clathrates, however, have significant contributions from interband transport that increases with increasing temperature. At 600 K, 8% and 11% of the total thermal conductivity for BaGG and SrGG, respectively, are due to interband transport. We note that our coarse-grained model neglects flat, high-frequency modes related to intra-cage dynamics [7], and contributions to interband transport may thus be more significant with an atomistic description of these clathrates [55].
@@ III.3 The role of nonperturbative effects on thermal conductivity
Figure 4: (a) Thermal conductivities of GG (blue), BaGG (green), and SrGG (red) calculated using the Green-Kubo formalism [Eq. (14)] with spectral functions calculated using SCP+VDMFT (triangles) and using the BTE formalism [Eq. (1)] with lifetimes calculated using SCP+PT (circles). All calculations are sampled on a $12\times 12\times 12$ Gamma-centered grid of the BZ. Dashed lines are a guide to the eye. (b) Thermal conductivities at higher temperatures, where black points are experimental measurements from Ref. [15]. Grey lines indicating different power laws are a guide to the eye.
@@
To better understand the role of nonperturbative effects on the thermal conductivity, we compare the thermal conductivity across materials and temperatures computed using two different methods: the Green-Kubo formalism [Eq. (14)] with spectral functions calculated using SCP+VDMFT (GK-VDMFT) and the BTE-SMA approach [Eq. (1)] with SCP phonons and lifetimes calculated using conventional lowest-order PT of the three-phonon scattering processes (BTE-PT). Details regarding the BTE-PT calculations are provided in the SM [33].
@@
When calculated using GK-VDMFT, the thermal conductivity for all materials peaks around 25-50 K and then decreases with increasing temperature. As discussed above, the introduction of guest atoms decreases the thermal conductivity by more than one order of magnitude, and BaGG, which features weaker cage-guest anharmonicity, has a larger thermal conductivity than SrGG at all temperatures. The BTE-PT approach results in thermal conductivity values that are consistent with GK-VDMFT for the quasi-harmonic GG, but they disagree significantly for the filled clathrates. BTE-PT predicts only a factor of two reduction between GG and BaGG, and it incorrectly predicts that the more anharmonic SrGG has a larger thermal conductivity than BaGG. These discrepancies are a result of the failure of SCP+PT to capture the short lifetimes of the filled clathrates (Fig. 1d), causing a significant overestimation of their thermal conductivities. As shown in the SM [33], using the BTE-SMA formalism with lifetimes determined from SCP+VDMFT calculations yields values that closely align with those calculated by the GK-VDMFT approach, suggesting that the success of the BTE-SMA depends strongly on the quality of the phonon lifetimes used and, in this case, is not due to the exclusion of interband transport. The SM [33] also illustrates comparisons to thermal conductivities calculated in the classical limit [Eq. (11)] using molecular dynamics simulations, showing consistency with the GK-VDMFT results. These comparisons suggest that the ultralow thermal conductivities of the filled clathrates result from four-phonon and higher-order scattering processes that are not described by SCP+PT but that are inherently included in our SCP+VDMFT approach, highlighting the importance of nonperturbative anharmonic effects on observables like the thermal conductivity. In Fig. 4b, we compare our calculated values of the thermal conductivity for BaGG to experimental data, focusing on high-temperature data, where intrinsic effects of anharmonicity dominate over effects due to grain boundary scattering, dynamical disorder, and nuclear quantum effects. We see excellent agreement although the degree of quantitative agreement needs to be further understood in light of the approximations made in our coarse-grained model. At lower temperatures, phonon-grain boundary scattering can become important [7], as discussed in the SM [33].
@@
Finally, we consider the temperature dependence of the total thermal conductivity (Fig. 4b), as strong interactions can produce anomalous behavior, as in electronic transport properties [56]. Here, we focus on the high-temperature dependence, as our calculations neglect nuclear quantum effects and grain boundary scattering, which may affect the behavior of thermal conductivities at low temperatures. For all materials studied here, BTE-PT predicts a $T^{-1}$ temperature dependence of the thermal conductivity. This temperature dependence is dictated by that of the phonon lifetimes, for which a $T^{-1}$ scaling is inherent to the PT formalism at temperatures greater than the phonon frequency (see SM [33]). In contrast, the GK-VDMFT approach predicts different high-temperature behavior for the different clathrates. BaGG has an intermediate scaling of $\sim T^{-0.9}$, and SrGG, the most anharmonic system, has the most moderate scaling, $\sim T^{-0.6}$. While interband transport can account for deviations from the $T^{-1}$ scaling in some materials [21], the good agreement between the GK-VDMFT and BTE-VDMFT approaches (illustrated in the SM [33]) suggest that interband transport is not crucial to this behavior in the clathrate models studied here.
@@
The high-temperature dependence of the thermal conductivity primarily arises from that of the acoustic phonon lifetimes. Thus, we analyze the temperature dependence of longitudinal and transverse acoustic phonon modes near the BZ center of BaGG and SrGG, as illustrated in Fig. 5. While the mode lifetimes of BaGG largely follow a $T^{-1}$ behavior, the mode lifetimes of SrGG deviate significantly from the $T^{-1}$ behavior that is predicted by perturbation theory and begin to saturate to a milder scaling at 50 K. This high-temperature saturation of the phonon lifetimes, and thus thermal conductivities, is reminiscent of “resistivity saturation” in metals [56, 57, 58, 59] and is a hallmark of nonperturbative interactions. This result indicates that the acoustic modes of the SrGG cage framework are subject to multiphonon scattering processes that are not captured by conventional PT. Despite their relatively narrow linewidths and long lifetimes (Fig. 1), these acoustic phonons’ low energies and strong coupling to Sr(2) rattling modes makes them more likely to participate in multiphonon scattering processes. Thus, we find that including nonperturbative descriptions of anharmonicity is essential in accurately capturing thermal transport properties, especially at higher temperatures where multiphonon processes are activated.